### v1

@@ Abstract
The performance of quantum key distribution (QKD) heavily depends on the finite statistics of its security proof. For multiple protocols and proof techniques, the central statistical task is a random sampling problem, which is customarily addressed by invoking suitable tail bounds on the hypergeometric distribution. In this work, we introduce an alternative solution that exploits a link between random sampling with and without replacement. Despite its simplicity, it notably boosts the achievable secret key rate, particularly in the regime of small data block sizes critical for satellite QKD and other envisioned QKD applications. Moreover, as a by-product of the proposed tool, tight Neyman constructions are derived for the average of independent Bernoulli variables. Bounds of this kind naturally fit in finite-key security proofs of decoy-state QKD schemes, further sharpening the finite statistics compared to previous approaches.
Introduction. Quantum key distribution (QKD) is the only approach to key exchange that is provably secure against computationally unbounded adversaries [1]. Today, considerable efforts are being made to boost the performance of practical QKD systems while guaranteeing their physical-layer security [2, 3]. Among the multiple facets of this ongoing program, refining the mathematical security proofs is of utmost importance in order to achieve tighter lower bounds on the extractable secret key length. These security proofs build upon quantum information-theoretic results and rely on statistical prediction and inference. Here, finite statistics play a pivotal role to keep the necessary data block sizes to a minimum. This feature is very desirable for any application of QKD and may be critical for space-to-ground QKD links in low or medium Earth-orbit environments [4, 5, 6, 7, 8], where both the raw data collection per satellite overpass and the data storage capacity of the satellite are limited [10, 11, 9].
Establishing the finite-key security of a QKD protocol requires to upper bound the probability of a “failure” happening in its parameter estimation (PE) step. This refers to the event where the protocol does not abort but the adversary obtains non-trivial information about the distributed key [1]. In BB84-like protocols [12], a very popular security proof technique is based on the entropic uncertainty relation [13], in which case the simplest formulation of the problem is as follows. Let us consider a two-piece experiment where a population of $N$ binary random variables (RVs) is generated according to an arbitrary probability distribution, and a random sample of size $n$ is drawn from the population afterwards. If we denote the frequency of ones in the test (complementary) sample as $\hat{p}$ ($\hat{q}$), the goal is to find a threshold value $q^{\rm th}\geq{}p^{\rm th}$ such that
@@
for arbitrary values of $N$, $n$, $\epsilon$ and $p^{\rm th}$. Naturally, an interplay exists between $\epsilon$ and $q^{\rm th}$: fulfilling Eq. (1) with a very small $\epsilon$ comes at the price of tolerating larger values of $q^{\rm th}$. In the context of QKD —in which $\hat{q}$ is often referred to as the phase-error rate (PHER)— this translates into a direct penalty on the provably secure key length. Remarkably, the above formulation of the problem shows up exactly in ideal implementations of the BB84 [12] or the BBM92 [14, 15] protocols, and we adhere to the latter in the main text for didactic purposes. See Appendix A for the detailed protocol description and secret key length formula. On the other hand, the popular decoy-state BB84 protocol [16, 17, 18, 19], which is implemented in most of today’s commercial QKD setups, is addressed in Appendix B.
A threshold value $q^{\rm th}$ satisfying Eq. (1) can be obtained from an exponential tail bound on the hypergeometric distribution known as Serfling inequality [20], an approach that is commonplace in QKD security proofs nowadays [15, 21]. Recently, though, a sharper bound has been derived in [22] by suitably splitting the “failure event” into two disjoint scenarios delimited by the total frequency of ones in the population, and finding separate estimates for each of them. In particular, while the first estimate still relies on the original Serfling bound, the second one is based on a refined inequality due to Hush and Scovel [23], resulting in a 14% to 17% reduction of the minimum data block length necessary for the BBM92 scheme with typical channel and protocol settings. Notably though, the improvement reported in [22] with respect to [15] heavily relies on the use of considerably large samples for testing. This is an undesirable feature because the final key is extracted from the complementary data.
In this paper, we propose a fundamentally different approach to calculate $q^{\rm th}$, which exploits a link between random sampling with and without replacement [24]. Remarkably, the resulting bound outperforms [22] in all the explored parameter regimes relevant for QKD, where one naturally restricts $p^{\rm th}$ to a few percent at most, $N$ and $n$ to thousands or tens of thousands at least, and $\epsilon$ to rather minuscule values. In particular, a significant improvement is observed for small data block sizes, where, to the best of our knowledge, [22] provides the tightest analysis in the literature. Precisely, for the BBM92 protocol, we find reductions of the minimum data block length beyond 16% in practically relevant scenarios, while for the decoy-state BB84 scheme we report a reduction of 22% when compared to [22], which is remarkable. Moreover, the advantage becomes larger when considering more demanding (i.e. smaller) values of the failure probability $\epsilon$. On another note, our bound also surpasses that of [22] in a vast parameter regime alien to QKD, which suggests that it might be of independent interest to other applications.
Furthermore, in the process of deriving the main result, tight analytical Neyman constructions [25] are obtained for the average of independent Bernoulli variables. Importantly, these naturally fit in e.g. finite-key security analyses of decoy-state QKD schemes, thereby allowing to further sharpen the finite statistics.
Technical results. The key enabler of our method is the elementary observation that tail bounds for binary samples drawn with replacement also hold for samples drawn without replacement, which is a consequence of a more general theorem by Hoeffding [24]. The derivation of the main result involves two conceptual steps. First, Neyman constructions [25] —essentially, confidence intervals— are derived from the additive Chernoff bound [26], a well-known exponential tail bound for sampling with replacement. This step relies on a tight-as-possible rational approximation of the logarithmic function [27] to provide a relaxation of the additive Chernoff bound. Secondly, the upper Neyman construction, which directly extends to the hypergeometric distribution, is recasted in terms of a hypergeometric RV and its complementary RV, and the desired bound follows through an averaging technique similar to those of [21, 15, 22].
First, we provide the statement of the Neyman constructions, a proof of which is given in Appendix C.
Proposition 1 (Neyman constructions).
@@
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, with expected value $\mathbb{E}\left[\hat{p}\right]=p$, and let $\kappa_{n,\epsilon}=(2/9n)\ln(1/\epsilon)$. Then, for all $\epsilon>0$, we have that $\Pr[\Gamma^{+}_{n,\epsilon}(\hat{p})\leq{}p]\leq{\epsilon}$, where $\Gamma^{+}_{n,\epsilon}(x)$ is given by the upper branch of
@@
±$\pm 3\sqrt{\kappa_{n,\epsilon}\left(\kappa_{n,\epsilon}+x-x^{2}\right)}\biggr{]}$
for $x\leq{}(1-2\kappa_{n,\epsilon})/(1+\kappa_{n,\epsilon})$, while for $x>(1-2\kappa_{n,\epsilon})/(1+\kappa_{n,\epsilon})$ the function $\Gamma^{+}_{n,\epsilon}(x)$ is extended with the slope at $x=(1-2\kappa_{n,\epsilon})/(1+\kappa_{n,\epsilon})$.
Also, for all $\epsilon>0$, we have that $\Pr[\Gamma^{-}_{n,\epsilon}(\hat{p})\geq{}p]\leq{\epsilon}$, where $\Gamma^{-}_{n,\epsilon}(x)$ is given by the lower branch of Eq.(2) for $x\geq{}3\kappa_{n,\epsilon}/(1+\kappa_{n,\epsilon})$, while for $x<3\kappa_{n,\epsilon}/(1+\kappa_{n,\epsilon})$ the function $\Gamma^{-}_{n,\epsilon}(x)$ is extended with the slope at $x=3\kappa_{n,\epsilon}/(1+\kappa_{n,\epsilon})$.
These constructions automatically extend to a hypergeometric RV because the additive Chernoff bound —from which they are solely derived— also holds in that setting (see Theorem 4 in [24]). Particularly, the extension of the upper Neyman construction is given below.
@@
Let $\hat{X}\sim{}Hypergeometric(N,K,n)$ and $\hat{p}=\hat{X}/n$, such that $\mathbb{E}\left[\hat{p}\right]=p$ with $p=K/N$. Then, for all $\epsilon>0$, we have that $\Pr[\Gamma^{+}_{n,\epsilon}(\hat{p})\leq{}p]\leq{\epsilon}$, where $\Gamma^{+}_{n,\epsilon}(x)$ is defined in Eq. (2).
@@
Proposition 2 (Random sampling bound).
A test sample of size $n$ is drawn at random from a binary population with $N$ elements. Let $\hat{p}$ ($\hat{q}$) denote the frequency of ones in the test (complementary) sample. For all $\epsilon>0$, we have that
$\Pr\left[\hat{q}\geq{}q^{\rm th}_{N,n,\epsilon}(\hat{p})\right]\leq{}\epsilon,$ (3)
@@
where $\Gamma^{+}_{n,\epsilon}(x)$ is defined in Eq. (2).
@@
Let $\hat{X}$ be the number of ones in the test sample, such that $\hat{p}=\hat{X}/n$. Since $\hat{X}$ is hypergeometrically distributed, the claim follows directly from Corollary 1, by substituting $p=n\hat{p}/N+(N-n)\hat{q}/N$ and rearranging terms. ∎
Importantly, as shown next, Proposition 2 allows to obtain a threshold value $q^{\rm th}$ fulfilling Eq. (1), thus directly providing the failure probability estimation for the ideal BBM92 protocol.
Proposition 3 (BBM92 failure probability).
Consider the two-piece experiment presented in the Introduction. Then, for all $\epsilon>0$ and $p^{\rm th}<1/2$, we have that
$\Pr[\hat{p}\leq{}p^{\rm th},\hat{q}\geq{}q^{\rm th}_{N,n,\epsilon}\left(p^{\rm th}\right)\Bigr{]}\leq{}\epsilon,$ (5)
for the function $q^{\rm th}_{N,n,\epsilon}(x)$ defined in Eq. (4), given that the latter is locally non-decreasing at $p^{\rm th}$.
@@
A relevant observation is that the condition on the slope at $p^{\rm th}$, which is equivalent to $\Gamma^{+^{\prime}}_{n,\epsilon}(p^{\rm th})\geq{}n/N$ ($\Gamma^{+^{\prime}}_{n,\epsilon}(x)$ denoting the derivative of $\Gamma^{+}_{n,\epsilon}(x)$ with respect to $x$), is unrestrictive for all practical purposes [28]. On another note, we remark that Proposition 2 also provides a building block for the estimation of the failure probability of other QKD schemes; see Appendix B.
In Fig. 1 we compare the threshold values $q^{\rm th}$ of Eq. (1) arising from Proposition 3, [22] and [15]. Precisely, we plot $q^{\rm th}$ versus $p^{\rm th}$ for a range $p^{\rm th}\in[0.01,0.05]$, setting $N=10^{6},n=N/4$, and $\epsilon=10^{-10}$ for illustration purposes. Notably, the improvement attained by Proposition 3 increases further for lower values of $p^{\rm th}$, $n$ and $\epsilon$.
Figure 1: Estimation of $q^{\rm th}$ versus $p^{\rm th}$ when $p^{\rm th}\in[0.01,0.05]$, $N=10^{6},n=N/4$, and $\epsilon=10^{-10}$. The solid blue line corresponds to Proposition 3 of this work (i.e. $q^{\rm th}_{N,n,\epsilon}\left(p^{\rm th}\right)$), while the dashed orange line refers to [22], which combines Serfling [20] and Hush and Scovel [23] inequalities. The dotted green line refers to [15]
, which is based on Serfling inequality alone.
QKD simulations. In what follows, we illustrate the impact that the proposed statistical tools have in the finite-key performance of QKD.
@@
We consider the ideal BBM92 protocol first. Consistently with [22, 15], here we evaluate a simplified scenario in which Alice and Bob share a $2N$-partite quantum state as protocol input, and they jointly select between two binary and complementary quantum measurements uniformly at random on a round-by-round basis, in so each of them extracting a raw data block of $N$ bits [15]. These blocks are randomly splitted into a fixed-length test (sifted-key) string of $n$ ($N-n$) bits. That is to say, both the key and the test string are randomly sampled from the two bases.
In the simulations, we set the tolerated quantum bit error rate (QBER) to $p^{\rm th}=4.55\%$ for ease of comparison with [22], where this choice is motivated by the results of the Micius experiment [6]. Also, an error correction (EC) leakage of $\lambda_{\rm EC}=1.19(N-n)h(p^{\rm th})$ bits is assumed, $h(x)$ standing for the binary Shannon entropy of $x$. The correctness parameter, $\epsilon_{\rm cor}$, and the privacy amplification error, $\epsilon_{\rm PA}$, are both set to $10^{-8}$ for illustration purposes. On the other hand, the failure probability bound is set to $\epsilon_{\rm PE}=4\times{}10^{-16}$, which roughly provides the minimum value enabling a low enough $q^{\rm th}$ for a nonzero secure key length in the Micius experiment. Altogether, these settings lead to an overall secrecy parameter of $\epsilon_{\rm sec}\approx{}5\times{}10^{-8}$ (see Appendix A).
With these inputs, in Fig. 2a we plot the extractable secret key length $l$ divided by the data block size for data blocks ranging from $N=2000$ to $8000$ bits, showing that our method allows to reduce the minimum block size with respect to [22] in more than 16%. As a side remark, we also note that the approach in [22] would require $\epsilon_{\rm PE}$ to be at least $4\times 10^{-13}$ to obtain a positive key in the Micius experiment, which is three orders of magnitude larger than what is needed with our tool. In addition, we emphasize that the advantage of our approach increases further if more demanding (i.e. smaller) values of $\epsilon_{\rm cor}$ and $\epsilon_{\rm PA}$ are requested.
(a) Ideal BBM92 protocol
(b) Decoy-state BB84 protocol
Figure 2: Finite secret key rate per raw data bit as a function of the data block size. (a) Ideal BBM92 protocol. The solid blue line corresponds to Proposition 3 of this work, while the dashed orange (dotted green) line refers to [22] ([15]). The dotted red vertical line indicates the block size gathered in the Micius BBM92 experiment [6], i.e. $N=3100$ bits. (b) Decoy-state BB84 protocol. The solid blue line corresponds to Propositions 1 and 2 of this work, while the dashed orange (dotted green) line refers to the combination of [22] with [30] ([24]). The experimental parameters considered in the simulations are provided in the text.
Let us now consider the more practical decoy-state BB84 protocol. We assume a scheme with three common intensities per basis, $\mu$ (signal), $\nu$ (decoy) and $\omega$ (vacuum), and deploy the $Z$ ($X$) basis for key extraction (PE) with a fixed-size data block of $N_{\rm Z}$ ($N_{\rm X}$) bits drawn at random from the corresponding data pool. In the PE step of the protocol, the parties evaluate suitable decoy-state bounds for the single-photon security parameters (e.g. the number of single-photon counts and their PHER). If these bounds do not fulfil predetermined acceptance thresholds, the protocol aborts. Otherwise, a fixed-length output key matching the acceptance thresholds is extracted. As explained in Appendix B, the details of the PE test are different depending on the random sampling technique used to bound the PHER.

### v2

@@ Abstract
The performance of quantum key distribution (QKD) heavily depends on statistical inference. For a broad class of protocols, the central statistical task is a random sampling problem, customarily addressed using exponential tail bounds on the hypergeometric distribution. Here we devise a strikingly simple exponential bound for this task, of unprecedented tightness among QKD security analyses. As a by-product, confidence intervals for the average of non-identical Bernoulli parameters follow too. These naturally fit in statistical analyses of decoy-state QKD and also outperform standard tools. Lastly, we show that, in a vast parameter regime, the use of tail bounds is not enforced because the cumulative mass function of the hypergeometric distribution is accurately computable. This sharply decreases the minimum block sizes necessary for QKD, and reveals the tightness of our simple analytical bounds when moderate-to-large blocks are considered.
Introduction. Quantum key distribution (QKD) is the only approach to key exchange that is provably secure against computationally unbounded adversaries [1]. Today, considerable efforts are being made to boost the performance of practical QKD systems while guaranteeing their physical-layer security [2, 3]. Among the multiple facets of this ongoing program, it is of utmost importance to achieve tighter lower bounds on the extractable secret key length, a goal that can be tackled by refining QKD security proofs. These proofs build upon quantum information-theoretic results and strongly rely on statistical inference, which plays a pivotal role to keep the necessary data block sizes to a minimum.
To be precise, establishing the finite-key security of a QKD protocol requires to upper bound the probability of a “failure” happening in its parameter estimation (PE) step. This refers to the event where the protocol does not abort but the adversary obtains non-trivial information about the final key [1]. In BB84-like protocols [4], a very popular proof technique is based on an entropic uncertainty relation (EUR) for smooth entropies [5, 6], in which case the simplest formulation of the problem goes as follows. Let us consider a two-piece experiment where a population of $N$ binary random variables is generated according to an arbitrary probability distribution, and a random sample of size $n$ is drawn from the population afterwards. If we denote the frequency of ones in the test (complementary) sample as $\hat{p}$ ($\hat{q}$), the goal is to find a threshold value $q^{\rm th}\geq{}p^{\rm th}$ such that
@@
for arbitrary values of $N$, $n$, $\epsilon$ and $p^{\rm th}$. Naturally, an interplay exists between $\epsilon$ and $q^{\rm th}$: fulfilling Eq. (1) with a very small $\epsilon$ comes at the price of tolerating larger values of $q^{\rm th}$. In the context of QKD —in which $\hat{q}$ is often referred to as the phase-error rate (PHER)— this translates into a direct penalty on the provably secure key length. Notably, the above formulation of the problem shows up exactly in ideal implementations of the BB84 [4] or the BBM92 [7, 8] protocols, and we adhere to the latter in the main text for didactic purposes. The detailed protocol description and the secret key length formula can be found in Supplemental Material I. On the other hand, the popular decoy-state BB84 protocol [9, 10, 11, 12], which is implemented in most of today’s commercial QKD setups, is addressed in Supplemental Material II.
In EUR-based security proofs [8, 13], it is commonplace to derive the desired $q^{\rm th}$ by resorting to an exponential tail bound on the hypergeometric (HG) distribution known as Serfling inequality [14]. Recently though, a sharper bound has been derived in [15], combining Serfling inequality with a refined inequality due to Hush and Scovel [16]. Nonetheless, the attainable improvement is moderate and heavily relies on the use of very large samples for testing, an undesirable feature given the fact that the final key is extracted from the complementary data.
In this paper, we propose a simpler approach to calculate $q^{\rm th}$, resorting to a well-known link between random sampling with and without replacement [17]. Remarkably, the bound we obtain outperforms [15] in all the explored parameter regimes relevant for QKD. Furthermore, in the derivation of the bound, confidence intervals (CIs) are obtained for the average parameter of a set of independent Bernoulli trials. These intervals, which provide an additional tool required for current decoy-state QKD security proofs, also outperform the analytical results considered for that purpose so far [12, 18]. Lastly, we show that, contrary to general belief, the cumulative mass function (CMF) of the HG distribution is exactly computable in a vast parameter regime of interest to QKD. Crucially, this enables the calculation of optimal —so-called Clopper-Pearson— one-sided CIs for the population parameter [19]. Applying the latter to the random sampling problem of Eq. (1) —and also to its generalization to decoy-state QKD schemes— we report a dramatic decrease of the minimum block sizes necessary for key extraction. On the other hand, explicit comparison with our simple analytical tools reveals the tightness of the latter for moderate-to-large blocks.
Analytical results. Our analytical results are triggered by two observations. First, the additive Chernoff bound [20], conventionally used in the context of sampling with replacement, also holds for sampling without replacement. This feature, which is a consequence of a more general theorem by Hoeffding [17] —see also [21] for a simple proof— has already been exploited in QKD security analyses before [22, 23]. Secondly, regardless of the context, very loose relaxations of the bound are often used in the literature. With this in mind, we derive a tight relaxation of the additive Chernoff bound and indistinctly use it to obtain one-sided CIs for the population parameters in sampling with and without replacement. Then, we apply these CIs to address the problem of Eq. (1) and its generalization to decoy-state QKD.
Crucially, our relaxation stems from a rational approximation of the Kullback-Leibler divergence, which is based on a logarithmic inequality [24] carefully selected to keep the CIs tight and handy at the same time. Indeed, higher-order rational approximations of the logarithm lead to much more convoluted formulas for the divergence. The reader is referred to Supplemental Material III for the relaxed Chernoff bound, and the corresponding CIs are stated in Proposition 1 and Corollary 1 below.
Proposition 1.
@@
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, with expected value $\mathbb{E}\left[\hat{p}\right]=p$, and let $\kappa_{n,\epsilon}=(2/9n)\ln(1/\epsilon)$. Then, for all $\epsilon>0$, $\Pr[\Gamma^{-}_{n,\epsilon}(\hat{p})\geq{}p]\leq{\epsilon}$ and $\Pr[\Gamma^{+}_{n,\epsilon}(\hat{p})\leq{}p]\leq{\epsilon}$, where
$\Gamma^{-}_{n,\epsilon}(\hat{p})=\left\{\begin{array}[]{ll}\gamma^{-}_{n,\epsilon}(\hat{p})&\mathrm{if}\hskip 5.69046pt\hat{p}\in{}\left[\displaystyle{\frac{3\kappa_{n,\epsilon}}{1+\kappa_{n,\epsilon}}},1\right],\\ -\epsilon&\mathrm{otherwise},\end{array}\right.$ (2)
$\Gamma^{+}_{n,\epsilon}(\hat{p})=\left\{\begin{array}[]{ll}\gamma^{+}_{n,\epsilon}(\hat{p})&\mathrm{if}\hskip 5.69046pt\hat{p}\in\left[0,\displaystyle{\frac{1-2\kappa_{n,\epsilon}}{1+\kappa_{n,\epsilon}}}\right],\\ 1+\epsilon&\mathrm{otherwise},\end{array}\right.$ (3)
and
@@
±$\pm 3\sqrt{\kappa_{n,\epsilon}\left(\kappa_{n,\epsilon}+x-x^{2}\right)}\biggr{]}.$
A proof of this claim is given in Supplemental Material III. Importantly, since the claim is solely derived from the additive Chernoff bound, it automatically extends to a HG random variable [17, 21]. The upper extension goes as follows.
@@
Let $\hat{X}\sim{}Hypergeometric(N,K,n)$ and $\hat{p}=\hat{X}/n$, such that $\mathbb{E}\left[\hat{p}\right]=p$ with $p=K/N$. Then, for all $\epsilon>0$, $\Pr[\Gamma^{+}_{n,\epsilon}(\hat{p})\leq{}p]\leq{\epsilon}$, where $\Gamma^{+}_{n,\epsilon}(x)$ is defined in Proposition 1.
@@
Proposition 2.
A test sample of size $n$ is drawn at random from a binary population with $N$ elements. Let $\hat{p}$ ($\hat{q}$) denote the frequency of ones in the test (complementary) sample. For all $\epsilon>0$,
$\Pr\left[\hat{q}\geq{}q^{\rm th}_{N,n,\epsilon}(\hat{p})\right]\leq{}\epsilon$ (5)
@@
where $\Gamma^{+}_{n,\epsilon}(x)$ is defined in Proposition 1.
@@
Let $\hat{X}$ be the number of ones in the test sample, such that $\hat{p}=\hat{X}/n$. If we denote by $p$ the fraction of ones in the population, $\hat{X}\sim{}\mathrm{Hypergeometric}(N,Np,n)$ and the claim follows directly from Corollary 1 by substituting $p=n\hat{p}/N+(N-n)\hat{q}/N$ and rearranging terms. ∎
As shown next, Proposition 2 allows to obtain a threshold value $q^{\rm th}$ fulfilling Eq. (1), in so providing the failure probability estimation of the BBM92 protocol.
Proposition 3.
Consider the two-piece experiment presented in the Introduction. Then, for all $\epsilon>0$ and $p^{\rm th}<1/2$,
$\Pr[\hat{p}\leq{}p^{\rm th},\hat{q}\geq{}q^{\rm th}_{N,n,\epsilon}\left(p^{\rm th}\right)\Bigr{]}\leq{}\epsilon$ (7)
for the threshold function $q^{\rm th}_{N,n,\epsilon}(x)$ defined in Eq. (6), given that the latter is locally non-decreasing at $p^{\rm th}$.
@@
A relevant observation is that the condition on the slope at $p^{\rm th}$, which is equivalent to $\Gamma^{+^{\prime}}_{n,\epsilon}(p^{\rm th})\geq{}n/N$ ($\Gamma^{+^{\prime}}_{n,\epsilon}(p^{\rm th})$ denoting the derivative of $\Gamma^{+}_{n,\epsilon}(x)$ with respect to $x$ evaluated in $p^{\rm th}$), is unrestrictive for all practical purposes [25].
On another note, the reader is referred to Supplemental Material II for the failure probability estimation of a decoy-state QKD protocol, which relies not only on Proposition 2 but also on Proposition 1.
Numerical results. Certainly, tighter results can be obtained by resorting to numerical methods. For instance, using Stirling’s formula and the Gaussian cumulative density function, the authors of [22] construct a HG tail inequality that outpeforms our relaxed Chernoff bound in certain regimes, but does not enable a closed formula for the related CI. Alternatively, one could use the exact additive Chernoff bound rather than a relaxation. Note, however, that if one is willing to use numerical techniques, tail inequalities may become superfluous. This is so because, contrary to general belief, accurately computing the CMF of the HG distribution is often feasible in QKD applications. Crucially, this enables the calculation of the strictest possible one-sided CIs [26, 27] using the well-known Clopper-Pearson method based on pivoting the CMF [19]. In particular, for our purposes, one can replace $\Gamma^{+}_{n,\epsilon}(\hat{p})$ in Corollary 1 by the tighter statistic $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$, defined via [27]
$\mathcal{CP}^{+}_{N,n,\epsilon}\left(x\right)=\min\left\{p\geq{}x\big{\rvert}{}\Pr[\hat{p}\leq{}x|p\bigr{]}\leq{}\epsilon\right\}$ (9)
if the target set of the minimization is non-empty, and $\mathcal{CP}^{+}_{N,n,\epsilon}(x)=(N+1)/N$ otherwise [28], where we recall that $p$ takes discrete values in the set $\{0,1/N,2/N,\ldots{}1\}$ [29]. Remarkably, accurate determination of the CMF is required in order to compute $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$, with a numerical precision beyond $\epsilon$ according to Eq. (9). Indeed, for a broad range of settings of interest to QKD, this is a manageable task with existing software.
Importantly, since $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ fulfills Corollary 1, it also fulfills Proposition 2, and monotonicity considerations allow to further establish Proposition 3 replacing Eq. (6) by
$q^{\rm th}_{N,n,\epsilon}(x)=\frac{N\mathcal{CP}^{+}_{N,n,\epsilon}(x)-nx}{N-n}.$ (10)
In Fig. 1 we compare the candidate threshold values $q^{\rm th}$ satisfying Eq. (1): Eq. (6), Eq. (10), and the thresholds derived in [15] and [8]. For illustration purposes, we consider typical QKD inputs, $p^{\rm th}\leq{}4\%$, $N=10^{5}$, $n=10^{4}$ and $\epsilon=10^{-9}$, and remark that the improvement of our tools —i.e. Eq. (6) and Eq. (10)— prevails in all the explored parameter regimes relevant for QKD.
Figure 1: Threshold value $q^{\rm th}$ satisfying Eq. (1), for $p^{\rm th}\leq{}4\%$, $N=10^{5}$, $n=10^{4}$ and $\epsilon=10^{-9}$. The solid red (blue) line corresponds to our numerical (analytical) tool, based on the exact cumulative mass function of the hypergeometric distribution (a relaxed additive Chernoff bound [20]). The dashed orange line uses the result in [15], which combines Serfling [14] and Hush & Scovel [16] inequalities, and the dotted green line uses the result in [8], based on Serfling inequality alone.
On another note, in Supplemental Material IV we include optimal one-sided CIs for the average of independent Bernoulli parameters [30, 31], potentially replacing $\Gamma^{\pm}_{n,\epsilon}(\hat{p})$ (Proposition 1) in the failure probability estimation of decoy-state QKD schemes. These CIs are related to the well-known Clopper-Pearson CIs of the binomial distribution, which in fact have been used in the context of QKD security before (see e.g. [32, 33]).
@@
QKD simulations. Next, we show the impact that the proposed tools have in the finite-key performance of QKD. Let us consider the ideal BBM92 protocol first. Consistently with [15, 8], we evaluate a simplified scenario in which Alice and Bob share a $2N$-partite quantum state as a protocol input. Then, they jointly select between two binary and complementary quantum measurements uniformly at random on a round-by-round basis, in so each of them extracting a raw data block of $N$ bits. These blocks are randomly splitted into a fixed-length test (sifted-key) string of $n$ ($N-n$) bits. That is to say, both the key and the test string are randomly sampled from the two bases. In the simulations, we set the tolerated quantum bit error rate (QBER) to $p^{\rm th}=4.55\%$ for ease of comparison with [15], where this choice is motivated by the results of the Micius experiment [34]. Also, an error correction (EC) leakage of $\lambda_{\rm EC}=1.19(N-n)h(p^{\rm th})$ bits is assumed, $h(x)$ denoting the binary entropy of $x$. The correctness parameter, $\epsilon_{\rm cor}$, and the privacy amplification error, $\epsilon_{\rm PA}$, are both set to $10^{-8}$ for illustration purposes, and the failure probability bound is set to $\epsilon_{\rm PE}=4\times{}10^{-16}$. Altogether, these settings lead to an overall secrecy parameter of $\epsilon_{\rm sec}\approx{}5\times{}10^{-8}$ (see Supplemental Material I).
With these inputs, in Fig. 2a we plot the extractable key length divided by the data block size for blocks up to $N=10^{6}$ bits, optimizing the sampling fraction $n/N$. In addition, we superpose an inset of the minimum block-size regime, showing that our analytical tool (solid blue line) allows to reduce the minimum block size with respect to [15] (dashed orange line) by more than 16%, while our numerical tool (solid red line) attains a reduction beyond 40%. Furthermore, the tool of [15] would roughly require a three-orders-of-magnitude larger $\epsilon_{\rm PE}$ for a nonzero key with the block size collected in the Micius experiment [34], indicated with a dotted red vertical line in the inset. Notably as well, both of our tools perform very similarly —and significantly above [15, 8]— for non-critical block sizes.
Figure 2: Finite secret key rate versus data block size. (a) Ideal BBM92 protocol. The solid red (blue) line corresponds to our numerical (analytical) tool, the dashed orange line uses the result in [15], and the dotted green line uses the result in [8]. The dotted red vertical line in the inset indicates the block size gathered in the Micius experiment [34], $N=3100$ bits. (b) Decoy-state BB84 protocol. The same color code is used again, indicating the random sampling tool under consideration (in particular, the approach of [8] is not presented in this case). The additional confidence intervals required for the decoy-state analysis are selected as specified in the main text. For both figures, the settings employed in the simulations are provided in the main text as well.